OpenStudy (anonymous):

Simplify (m^3/3n)/^(-m^4 /9n2)

OpenStudy (anonymous):

this questions makes no sense

OpenStudy (anonymous):

^ you are missing something

OpenStudy (anonymous):

Do you mean (m^3/3n)/(-m^4/9n^2) ?

OpenStudy (anonymous):

(m^3/3n)/(-m^4 /9n^2)

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

\[\frac{(m^3/3n)}{(-m^4 /9n^2)}\]

OpenStudy (anonymous):

\[\frac{(a/b)}{(c/d)}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

See how we can rewrite this?

OpenStudy (anonymous):

so the answer be 3/mn

OpenStudy (anonymous):

-3n/m

OpenStudy (anonymous):

\[\frac{m^3}{3n} \frac{9n^2}{-m^4} = -\frac{3n}{m}\]

OpenStudy (anonymous):

can u help me on this one (6y^2 -6/ 8y^2 +8y)/(3y -3 / 4y^2 + 4y)

OpenStudy (anonymous):

Post a new question